# SeRL: Self-Play Reinforcement Learning for Large Language Models with Limited Data

### v1

@@ 1 Introduction
Recent breakthroughs of Large Language Models (LLMs) have demonstrated their potential to revolutionize a wide range of fields, including healthcare (Qiu et al., 2024; Li et al., 2025; Clusmann et al., 2023), robotics (Wang et al., 2024a; Kannan et al., 2024; Liu et al., 2024), and web services (Xiong et al., 2024b; Wu et al., 2024a; Gu et al., 2024). Notably, frontier models such as OpenAI-o3 (OpenAI, 2025) and DeepSeek-R1 (DeepSeek-AI et al., 2025a) have proven that Reinforcement Learning with Verifiable Rewards (RLVR) can endow LLMs with advanced reasoning capabilities (Xiong et al., 2024a; Wu et al., 2024a; Gu et al., 2024; Zhang et al., 2025a; Liu et al., 2025b), enabling significant achievements in mathematics (Yang et al., 2024; DeepSeek-AI et al., 2025a) and programming (Hui et al., 2024; Khan et al., 2025). A key factor in these developments is the availability of large-scale supervised datasets (Qwen et al., 2025; DeepSeek-AI et al., 2025b; Grattafiori et al., 2024; Yao et al., 2024), which incorporate instructions and corresponding ground-truth labels for reward computation.
However, in specialized domains such as clinical diagnostics (Ullah et al., 2024; McDuff et al., 2025; Wang and Liu, 2025) and aerospace engineering (Liu et al., 2025a; Connolly, 2025; Yadav, 2024), collecting such high-quality data and verifiable supervision remains a significant challenge. This difficulty stems from the fact that acquiring human-labeled data in these areas is particularly labor-intensive, often requiring substantial expert knowledge and considerable time investment (Silver and Sutton, 2025; Villalobos et al., 2024). Moreover, such human-generated annotations are inherently difficult to scale. As an alternative, data generated by LLMs offers a more scalable and cost-effective solution (Yang et al., 2023; Yuan et al., 2025b; Goldie et al., 2025; Ge et al., 2025; Li et al., 2024a). Nevertheless, this approach usually depends on access to highly capable expert-level LLMs, which may not always be readily available.
To remedy this issue, recent efforts have explored the self-instruction paradigm (Wang et al., 2023; Gao et al., 2024; Kim et al., 2025), where LLMs autonomously generate instructions and corresponding responses. While effective for supervised fine-tuning (Zhao et al., 2024; GALLOY et al., 2025; Guo et al., 2024), these methods are not directly applicable to online RL settings, which pose two critical challenges. First, existing methods typically generate a fixed dataset without accounting for the evolving capabilities of the LLM during online learning, rendering the static data progressively less suitable. Second, directly treating model-generated responses as ground-truth labels often introduces noise or inconsistency, leading to unreliable reward signals for RL.
@@
  • We explore how to leverage RL to incentivize the reasoning abilities of LLMs in data-scarce scenarios, a highly practical yet underexplored challenge for LLM research community.
@@ 2 Background
Reinforcement Learning for LLM. We denote the LLM parameterized by $\theta$ as $\pi_{\theta}$. For language generation, given an instruction $\boldsymbol{x}$, the model outputs a sequence $\boldsymbol{y}=(y_{1},\ldots,y_{T})$, where $y_{t}=\pi_{\theta}(\boldsymbol{x},\boldsymbol{y}_{<t})$ and $\boldsymbol{y}_{<t}=(y_{0},\ldots,y_{t-1})$. RL has recently proven effective in improving the logical reasoning abilities of LLMs. In the RL setting, we additionally require a reward signal. We denote the reward for a given input-output pair $(\boldsymbol{x},\boldsymbol{y})$ as $R(\boldsymbol{x},\boldsymbol{y})$. In this work, while our framework is not tied to any specific RL algorithm, we choose Reinforce++ (Hu et al., 2025) for its robustness and stable performance across diverse tasks. The RL training objective is defined as follows:
$\begin{split}\mathcal{L}_{\text{RL}}(\theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D},\boldsymbol{y}\sim\pi_{\theta}}\left[\frac{1}{|y|}\sum_{t=1}^{|y|}\min\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}\hat{A}(s_{t},y_{t}),\right.\right.\\ \left.\left.\operatorname{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\hat{A}(s_{t},y_{t})\right)\right],\end{split}$ (1)
where $\mathcal{D}$ is the dataset of instructions, $s_{t}=(\boldsymbol{x},\boldsymbol{y}_{<t})$, $\epsilon$ is a clipping hyperparameter introduced in PPO (Schulman et al., 2017) to stabilize training, and $\pi_{\theta_{\text{old}}}$ denotes the model parameters before the most recent update. Different RL algorithms compute the advantage term $\hat{A}(s_{t},y_{t})$ in different ways. In Reinforce++, the return is defined as:
$G(s_{t},y_{t})=\mathbb{I}(y_{t}=[\text{EOS}])R(\boldsymbol{x},\boldsymbol{y})-\beta\cdot\sum_{i=t}^{T}\frac{\pi_{\theta}(y_{i}\mid\boldsymbol{x},\boldsymbol{y}_{<i})}{\pi_{\text{ref}}(y_{i}\mid\boldsymbol{x},\boldsymbol{y}_{<i})},$ (2)
where the first term assigns the final reward when $y_{t}$ is the end-of-sequence token, while the second adds a per-token KL penalty (Miller et al., 2023). The computed returns are standardized across the entire training batch to compute the advantage estimate $\hat{A}(s_{t},y_{t})$.
@@
Figure 1: An overview of the proposed SeRL framework, which comprises two core components: (1) Self-Instruction, where the model generates new instructions from a small initial dataset and applies a robust online filtering strategy to ensure instruction quality, diversity, and appropriate difficulty. (2) Self-Rewarding, where the model performs unsupervised RL training using a majority-vote reward mechanism without relying on verifiable labels.
@@ 3.1 Self-Instruction
Instruction Generation. The LLM is prompted in a few-shot manner to generate new instructions. Each few-shot prompt consists of randomly selected examples from the initial dataset and examples from the already generated dataset. The reason we do not rely solely on the initial dataset as few-shot examples is that, as training progresses, we aim to adapt the examples to the evolving capabilities of the model. Specifically, we also selectively incorporate recently generated instructions as few-shot examples for subsequent generation, enabling the model to produce instructions that better match its current level of competence. The prompt template used here is provided in Tab. C.1.
Online Instruction Filter. To ensure the quality, diversity, and moderate difficulty of the generated instructions, we implement a robust online filtering strategy. Specifically, we discard instructions that meet any of the following criteria: (1) a ROUGE-L (Lin, 2004) score exceeding a predefined threshold with existing instructions, to prevent generating semantically similar instructions that reduce diversity; (2) the presence of specific keywords such as “image”, “graph”, or “picture”, which refer to visual content that LLMs cannot process; (3) excessively long or short, since long instructions often contain repetitive phrasing while short ones tend to lack neccesary context, both leading to invalid instructions; (4) having a majority answer ratio outside a specified range, ensuring that the retained instructions maintains suitable difficulty for training. Formally, we retain only instructions whose average reward $r_{\text{mean}}$ falls within two tunable bounds, $\gamma_{\text{diff}}\leq r_{\text{mean}}\leq\gamma_{\text{easy}}$.
The motivation behind this dual-end clipped difficulty filtering strategy is to avoid scenarios where the responses are either entirely correct without any information gain or entirely incorrect with no feasibility for improvement. Such extreme cases result in zero advantage in algorithms like Reinforce++ (Hu et al., 2025) and GRPO (Shao et al., 2024), which can lead to gradient vanishing and consequently significant model degradation. In the ablation studies presented in Section 4.3, we demonstrate that removing the difficulty filtering leads to reward hacking.
@@ 3.2 Self-Rewarding
In our self-instruction setup, only instructions are generated, and the absence of verifiable labels presents a key challenge for reward computation. Interestingly, Yue et al. (2025) suggests that RL with verifiable labels can be viewed as converting the Pass@K performance of a model into Pass@1. Inspired by this insight, we explore whether a similar benefit can be obtained in the absence of verifiable labels. To achieve this, as illustrated in the green region of Fig. 1, we introduce a majority-vote-based self-rewarding that serves two purposes: (1) it improves inference efficiency by enabling the model to reach Maj@N performance with just a single response (Pass@1) after RL training, (2) and it offers a straightforward yet effective way to estimate rewards without the need for verifiable labels. For a given instruction $\boldsymbol{x}_{i}$, we sample $n_{\text{vote}}$ responses, each consisting of a chain-of-thought reasoning $\boldsymbol{c}_{i}^{k}$ and a final answer $\boldsymbol{a}_{i}^{k}$:
@@
Since we adopt a majority-vote-based reward estimation method, we define the majority answer for a given instruction $\boldsymbol{x}_{i}$ as $\texttt{Maj}\left(\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{\text{vote}}}\right)$, where $\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{\text{vote}}}$ denotes the set of $n_{\text{vote}}$ answers sampled for $\boldsymbol{x}_{i}$. For each $\boldsymbol{a}_{i}^{k}$, we compute the corresponding reward $R_{i}^{k}=\texttt{Verify}\left(\boldsymbol{a}_{i}^{k},\ \texttt{Maj}\left(\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{vote}}\right)\right)$. The Verify function is defined as:
@@
As instances with all-correct or all-incorrect responses have already been eliminated by the online filter, a majority answer can always be identified. If there is a tie among multiple majority answers, the one with the shortest length is selected, based on the assumption that the shortest response is typically the most concise and correct. We refer to this self-rewarding strategy as majority-vote-based reward. It serves as a replacement for traditional reward computation methods that rely on reward models or verifiable labels, providing a practical and efficient approach to reward estimation.
@@ 3.3 Overall Framework
To enable continual improvement in data-scarce scenarios, we adopt a combination of self-instruction and self-rewarding. As shown in the yellow region of Fig. 1, the model iteratively generates new instructions from limited data and estimates rewards for sampled responses using majority voting. This process enables sustained reinforcement learning across training iterations. To match the evolving capability of the model, we generate a batch of instructions at each training step $t$ using self-instruction. For each instruction in $\mathcal{B}^{t}_{\text{gen}}$, we sample $n_{vote}$ responses and compute the majority-vote-based reward. The resulting triplets $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i},r_{i})\}$, consisting of the instruction, sampled response, and computed reward, are used to update the model parameters in one training step. At the next training step $t+1$, we repeat the same process to generate $\mathcal{B}^{t+1}_{\text{gen}}=\{\boldsymbol{x}_{i}\}^{n_{\text{rbs}}}$, followed by one step of RL training. This cycle repeats across iterations, gradually refining the performance of the model.
Table 1: Pass@1 comparison across benchmarks between our method and baseline approaches, with all models evaluated using greedy decoding. Initial refers to the original LLaMA-3.2-3B-Instruct or Qwen-2.5-7B-Instruct model. Bold indicates the best performance of each method across iterations, and underlined marks the best performance among all methods and all iterations.
@@
Initial 32.1 26.7 41.9 34.3 34.0 32.1 26.7 41.9 34.3 34.0
RL-GT 34.6 27.6 44.1 36.0 36.1 34.6 27.6 44.1 36.0 36.1
SeRL (iter1) 32.6 27.8 41.5 34.0 34.3 32.6 27.8 41.5 34.0 34.3
SeRL (iter2) 33.8 27.7 41.9 33.9 35.0 33.8 27.7 41.9 33.9 35.0
SeRL (iter3) 33.9 28.1 41.9 34.1 35.1 33.9 28.1 41.9 34.1 35.1
4 Experiments
4.1 Experimental Setup
Models and training settings. We conducted experiments using two model series: LLaMA-3.2-3B-Instruct (Meta-AI, 2025) and Qwen-2.5-7B-Instruct (Qwen et al., 2025). We simulate a data-scarce scenario in the mathematics domain by limiting the initial dataset to 500 instructions, which are uniformly sampled across all difficulty levels from the MATH training set (Hendrycks et al., 2021). We denote this initial set of 500 instructions as $\mathcal{D}_{\text{seed}}$. For training, we adopt the OpenRLHF (Hu et al., 2024) framework and employ its Reinforce++ (Hu et al., 2025) algorithm for reinforcement learning. Detailed training hyperparameters are provided in Tab. 5 and Tab. 6.
Baselines. We compare our method against two multi-round iterative training baselines and one RL baseline with ground-truth rewards (RL-GT): (1) Iterative RPO (I-RPO) (Pang et al., 2024), which runs for three iterations. In each iteration, four responses are sampled for each instruction in the MATH training set. Each response is scored using a rule-based reward, and the highest and lowest are selected as the chosen and rejected responses for DPO training. (2) Self-Rewarding DPO (SR-DPO) (Yuan et al., 2025a), also run for three iterations. In each iteration, a new dataset $\mathcal{D}_{\text{gen}}=\{\boldsymbol{x}_{i}\}^{N}_{i=1}$ is synthesized based on $\mathcal{D}_{\text{seed}}$ using the same self-instruction approach as ours. For each generated instruction, four responses are sampled, and model-based rewards are used to construct preference pairs for DPO training. (3) RL-GT is run for a single iteration, using the MATH training set for RL training. The rewards are computed using a rule-based reward function.
Since the original reward prompts in SR-DPO are intended for general tasks, we modify them to better fit mathematical scenarios, ensuring a fair comparison (see Tab. C.2). To ensure fairness in training data scale, we note that RL-GT and I-RPO use 7,500 instructions from the MATH training set, and SR-DPO generates 7,500 instructions per iteration. For our method, accounting for filtered data, we estimate that every 100 steps consumes roughly 7,500 instructions. Therefore, we regard every 100 steps of our method as one iteration to ensure a fair comparison with other baselines.
Evaluation Benchmarks. We evaluate model performance using five math-specific benchmarks and one general-purpose benchmark. (1) The math benchmarks include MATH-500 (Lightman et al., 2023), which focuses on medium-difficulty problems; MATH-Hard (Hendrycks et al., 2021), which contains level-5 competition problems from AMC and AIME to assess advanced reasoning; ASDiv (Miao et al., 2020), a dataset of 2,305 diverse elementary-level math word problems; College Math (Tang et al., 2024), which covers university-level topics such as algebra and calculus; and TabMWP (Lu et al., 2023), consisting of 38,431 problems that combine textual and tabular data to test multi-step reasoning. (2) To assess general reasoning ability, we include MMLU-Pro (Wang et al., 2024b), an enhanced version of MMLU covering STEM, humanities, and social sciences.

### v2

@@ 1 Introduction
Recent breakthroughs of Large Language Models (LLMs) have demonstrated their potential to revolutionize a wide range of fields [39, 90, 91], such as healthcare [4, 31, 51], robotics [25, 37, 61], and web services [16, 40, 67, 72]. Notably, frontier models such as OpenAI-o3 [49] and DeepSeek-R1 [6] have proven that Reinforcement Learning with Verifiable Rewards (RLVR) can endow LLMs with advanced reasoning capabilities [16, 38, 67, 71, 85, 86], enabling significant achievements in mathematics [6, 77] and programming [21, 26]. A key factor in these developments is the availability of large-scale supervised datasets [7, 14, 78, 80], which incorporate instructions and corresponding ground-truth labels for reward computation.
However, in specialized domains such as clinical diagnostics [44, 58, 60] and aerospace engineering [5, 36, 75], collecting such high-quality data and verifiable supervision remains a significant challenge. This difficulty stems from the fact that acquiring human-labeled data in these areas is particularly labor-intensive, often requiring substantial expert knowledge and considerable time investment [55, 59]. Moreover, such human-generated annotations are inherently difficult to scale. As an alternative, data generated by LLMs offers a more scalable and cost-effective solution [12, 13, 30, 79, 82]. Nevertheless, this approach usually depends on access to highly capable expert-level LLMs, which may not always be readily available.
To remedy this issue, recent efforts have explored the self-instruction paradigm [11, 27, 63], where LLMs autonomously generate instructions and corresponding responses. While effective for supervised fine-tuning [10, 17, 88], these methods are not directly applicable to online RL settings, which pose two critical challenges. First, existing methods typically generate a fixed dataset without accounting for the evolving capabilities of the LLM during online learning, rendering the static data progressively less suitable. Second, directly treating model-generated responses as ground-truth labels often introduces noise or inconsistency, leading to unreliable reward signals for RL.
@@
  • We explore how to leverage RL to incentivize the reasoning abilities of LLMs in data-scarce scenarios, a highly practical yet underexplored challenge for the LLM research community.
@@ 2 Background
Reinforcement Learning for LLMs. We denote the LLM parameterized by $\theta$ as $\pi_{\theta}$. For language generation, given an instruction $\boldsymbol{x}$, the model outputs a sequence $\boldsymbol{y}=(y_{1},\ldots,y_{T})$, where $y_{t}\sim\pi_{\theta}(\boldsymbol{x},\boldsymbol{y}_{<t})$ and $\boldsymbol{y}_{<t}=(y_{0},\ldots,y_{t-1})$. RL [22, 41, 56] has recently proven effective in improving the logical reasoning abilities of LLMs. In the RL setting, we additionally require a reward signal. We denote the reward for a given input-output pair $(\boldsymbol{x},\boldsymbol{y})$ as $R(\boldsymbol{x},\boldsymbol{y})$. In this work, while our framework is not tied to any specific RL algorithm, we choose Reinforce++ [19] for its robustness and stable performance across diverse tasks. The RL training objective is defined as follows:
$\begin{split}\mathcal{J}_{\text{RL}}(\theta)=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D},\boldsymbol{y}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{|y|}\sum_{t=1}^{|y|}\min\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}\hat{A}(s_{t},y_{t}),\right.\right.\\ \left.\left.\operatorname{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\hat{A}(s_{t},y_{t})\right)\right],\end{split}$ (1)
where $\mathcal{D}$ is the dataset of instructions, $s_{t}=(\boldsymbol{x},\boldsymbol{y}_{<t})$, $\epsilon$ is a clipping hyperparameter introduced in PPO [53] to stabilize training, and $\pi_{\theta_{\text{old}}}$ denotes the model parameters before the most recent update. Different RL algorithms compute the advantage term $\hat{A}(s_{t},y_{t})$ in different ways. The details of the Reinforce++ algorithm are provided in Appendix C.1.
@@
Figure 1: An overview of the proposed SeRL framework, which comprises two core components: (1) Self-Instruction, where the model generates new instructions from a small initial dataset and applies a robust online filtering strategy to ensure instruction quality, diversity, and appropriate difficulty. (2) Self-Rewarding, where the model performs unsupervised RL training using a majority-voting reward mechanism without relying on verifiable labels.
@@ 3.1 Self-Instruction
Instruction Generation. The LLM is prompted in a few-shot manner to generate new instructions. Each few-shot prompt consists of randomly selected examples from the initial dataset and examples from the already generated dataset. The reason why we use instructions from the initial dataset is that we aim to expand the dataset while maintaining the same distribution as the initial one, enabling the model to effectively learn the distribution of scarce data. The reason we do not rely solely on the initial dataset as few-shot examples is that, as training progresses, we aim to adapt the examples to the evolving capabilities of the model. Specifically, we also selectively incorporate recently generated instructions as few-shot examples for subsequent generation, enabling the model to produce instructions that better match its current level of competence. The prompt template of instruction generation is provided in Tab. C.4.
Online Instruction Filter. To ensure the quality, diversity, and moderate difficulty of the generated instructions, we implement a robust online filtering strategy. Specifically, we discard instructions that meet any of the following criteria: (1) a ROUGE-L [35] score exceeding a predefined threshold with existing instructions, to prevent generating semantically similar instructions that reduce diversity; (2) the presence of specific keywords such as “image”, “graph”, or “picture”, which refer to visual content that LLMs cannot process; (3) excessively long or short, since long instructions often contain repetitive phrasing while short ones tend to lack necessary context, both leading to invalid instructions; (4) having a majority answer ratio outside a specified range, ensuring that the retained instructions maintains suitable difficulty for training. Formally, we retain only instructions whose average reward $r_{\text{mean}}$ falls within two tunable bounds, $\gamma_{\text{diff}}\leq r_{\text{mean}}\leq\gamma_{\text{easy}}$. $r_{\text{mean}}$ is calculated as the majority answer ratio from multiple responses of the instruction in our majority-voting reward method.
The motivation behind this dual-end clipped difficulty filtering strategy is to avoid scenarios where the responses are either entirely correct without any information gain or entirely incorrect with no feasibility for improvement. Such extreme cases result in zero advantage in algorithms like Reinforce++ [19] and GRPO [54], which can lead to gradient vanishing [65, 42] and consequently significant model degradation. In the ablation studies presented in Section 4.3, we demonstrate that removing the difficulty filtering leads to reward hacking.
@@ 3.2 Self-Rewarding
In our self-instruction setup, only instructions are generated, and the absence of verifiable labels presents a key challenge for reward computation. Interestingly, Yue et al. [83] suggests that RL with verifiable labels can be viewed as converting the Pass@K performance of a model into Pass@1. Inspired by this insight, we explore whether a similar benefit can be obtained in the absence of verifiable labels. To achieve this, as illustrated in the green region of Fig. 1, we introduce a majority-voting self-rewarding that serves two purposes: (1) it improves inference efficiency by enabling the model to reach Maj@N performance with just a single response (Pass@1) after RL training, (2) and it offers a straightforward yet effective way to estimate rewards without the need for verifiable labels. For a given instruction $\boldsymbol{x}_{i}$, we sample $n_{\text{vote}}$ responses, each consisting of a chain-of-thought reasoning $\boldsymbol{c}_{i}^{k}$ and a final answer $\boldsymbol{a}_{i}^{k}$:
@@
Since we adopt a majority-voting reward estimation method, we define the majority answer for a given instruction $\boldsymbol{x}_{i}$ as $\texttt{Maj}\left(\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{\text{vote}}}\right)$, where $\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{\text{vote}}}$ denotes the set of $n_{\text{vote}}$ answers sampled for $\boldsymbol{x}_{i}$. For each $\boldsymbol{a}_{i}^{k}$, we compute the corresponding reward $R_{i}^{k}=\texttt{Verify}\left(\boldsymbol{a}_{i}^{k},\ \texttt{Maj}\left(\{\boldsymbol{a}_{i}^{k}\}_{k=1}^{n_{\text{vote}}}\right)\right)$. The Verify function is defined as:
@@
As instances with all-correct or all-incorrect responses have already been eliminated by the online filter, a majority answer can always be identified. If there is a tie among multiple majority answers, the one with the shortest length is selected based on the principle that a more concise response is preferred to reduce verbosity. We refer to this self-rewarding strategy as majority-voting reward. It serves as a replacement for traditional reward computation methods that rely on reward models or verifiable labels, providing a practical and efficient approach to reward estimation.
@@ 3.3 Overall Framework
To enable continual improvement in data-scarce scenarios, we adopt a combination of self-instruction and self-rewarding. As shown in the yellow region of Fig. 1, the model iteratively generates new instructions from limited data and estimates rewards for sampled responses using majority voting. This process enables sustained reinforcement learning across training iterations. To match the evolving capability of the model, we generate a batch of instructions at each training step $t$ using self-instruction. For each instruction in $\mathcal{B}^{t}_{\text{gen}}$, we sample $n_{vote}$ responses and compute the majority-voting reward. The resulting triplets $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i},r_{i})\}$, consisting of the instruction, sampled response, and computed reward, are used to update the model parameters in one training step. At the next training step $t+1$, we repeat the same process to generate $\mathcal{B}^{t+1}_{\text{gen}}=\{\boldsymbol{x}_{i}\}^{n_{\text{rbs}}}$, followed by one step of RL training. This cycle repeats across iterations, gradually refining the performance of the model.
Table 1: Pass@1 comparison across benchmarks between our method and baseline approaches, with all models evaluated using greedy decoding. Initial refers to the original LLaMA-3.2-3B-Instruct or Qwen-2.5-7B-Instruct model. Bold indicates the best performance.
@@
4 Experiments
4.1 Experimental Setup
Models and Training Settings. We conducted experiments using two model series: LLaMA-3.2-3B-Instruct [46] and Qwen-2.5-7B-Instruct [78]. We simulate a data-scarce scenario in the mathematics domain by limiting the initial dataset to 500 instructions, which are uniformly sampled across all difficulty levels from the MATH training set [18]. We denote this initial set of 500 instructions as $\mathcal{D}_{\text{seed}}$. For training, we adopt the OpenRLHF [20] framework and employ its Reinforce++ [19] algorithm for RL. Detailed training hyperparameters are provided in Tab. 6 and Tab. 7.
Baselines. We compare our method against two multi-round iterative training baselines and one RL baseline with ground-truth rewards (RL-GT): (1) Iterative RPO (I-RPO) [50], which runs for three iterations. In each iteration, four responses are sampled for each instruction in the MATH training set. Each response is scored using a rule-based reward, and the highest and lowest are selected as the chosen and rejected responses for DPO training. Rule-based methods [54, 50] determine correctness by extracting the answer from the response using regular expressions and comparing it against the ground-truth answer to assign a reward. (2) Self-Rewarding DPO (SR-DPO) [81], also run for three iterations. In each iteration, a new dataset $\mathcal{D}_{\text{gen}}=\{\boldsymbol{x}_{i}\}^{N}_{i=1}$ is synthesized based on $\mathcal{D}_{\text{seed}}$ using the same self-instruction approach as ours. For each generated instruction, four responses are sampled, and model-based rewards are used to construct preference pairs for DPO training. Model-based methods [81] rely on manually designed principles or scoring rubrics, which the model uses to evaluate the response and assign a reward. (3) RL-GT is run for a single iteration, using the MATH training set for RL training. The rewards are computed using a rule-based reward function.
Since the original reward prompts in SR-DPO are intended for general tasks, we modify them to better fit mathematical scenarios, ensuring a fair comparison (see Tab. C.5). To ensure fairness in training data scale, we note that RL-GT and I-RPO use 7,500 instructions from the MATH training set, and SR-DPO generates 7,500 instructions per iteration. For our method, we online-generate instructions during training, keeping the same number of training gradient steps as other baselines to ensure the total training data scale remains consistent, thereby guaranteeing a fair comparison.
Evaluation Benchmarks. We evaluate model performance using five math-specific benchmarks and one general-purpose benchmark. (1) The math benchmarks include MATH-500 [34], which focuses on medium-difficulty problems; MATH-Hard [18], which contains level-5 competition problems from AMC and AIME to assess advanced reasoning; ASDiv [47], a dataset of 2,305 diverse elementary-level math word problems; College Math [57], which covers university-level topics such as algebra and calculus; and TabMWP [43], consisting of 38,431 problems that combine textual and tabular data to test multi-step reasoning. (2) To assess general reasoning ability, we include MMLU-Pro [64], an enhanced version of MMLU covering STEM, humanities, and social sciences. All evaluations are conducted using greedy decoding with a maximum of 1,024 new tokens.
@@ 4.1 Experimental Setup
Initial 32.1 26.7 41.9 34.3 34.0 59.9 40.1 64.1 53.9 57.3
RL-GT 34.6 27.6 44.1 36.0 36.1 60.0 39.6 63.9 53.1 57.2
SeRL (iter1) 32.6 27.8 41.5 34.0 34.3 60.0 40.8 64.2 54.3 57.5
SeRL (iter2) 33.8 27.7 41.9 33.9 35.0 60.2 40.0 63.9 53.5 57.3
SeRL (iter3) 33.9 28.1 41.9 34.1 35.1 59.8 40.9 63.9 54.1 57.4